OpenStudy (anonymous):

Find the dot product, a • b. a = 6i + 5j, b = -5i + 4j

OpenStudy (anonymous):

whats so hard about this @Nairz 6i+5j*-5i+4j 6i(-5i+4j) +5(5i+4j) can you solve this?

OpenStudy (anonymous):

I haven't done this in a long time. Just need a refresher :)

OpenStudy (anonymous):

so are u refreshed now

OpenStudy (anonymous):

so just simplify that equation you wrote?

OpenStudy (cwrw238):

remember that i*j = 0

OpenStudy (anonymous):

yeah multiply it gonna show one 6i(-5i+4j) 6i*-5i + 6i*4j then simply it, do it for both and then add subract, pretty basic

OpenStudy (anonymous):

and when he says that i x j = 0 does that mean that 6i x 4j = 0? @snoo

OpenStudy (anonymous):

@cwrw238 pl0x cum

OpenStudy (cwrw238):

6i x 4j = 24*ij = 24*0 = 0

OpenStudy (anonymous):

^

OpenStudy (anonymous):

do same with other one, @Nairz

myininaya (myininaya):

doesn't <x,y> dot <c,d> mean xc+yd

OpenStudy (anonymous):

ok thanks @snoo and @cwrw238

OpenStudy (anonymous):

btw, call me daddy, i'm out :)

myininaya (myininaya):

\[<6,5> \cdot <-5,4>=6(-5)+5(4)\]

OpenStudy (anonymous):

omg yes.

OpenStudy (anonymous):

so -30 + 20 or 10

myininaya (myininaya):

well -30+20 is -10

OpenStudy (anonymous):

oh whoops

OpenStudy (anonymous):

i forgot to put in the - sign

OpenStudy (anonymous):

i'm an old man i m sorry i was thinking of multiplying as adding, im so sorry dear

OpenStudy (anonymous):

true @myininaya

myininaya (myininaya):

I think we all get things a little mixed up sometimes. Math is huge subject after all.

OpenStudy (anonymous):

thanks dear